Three resistors of 10, 12 and 15 ohms are connected in parallel. Evaluate the value of current to the parallel system that will make the current in the 10 ohm resistor equal to 2 A.


1
Expert's answer
2021-09-30T23:55:25-0400


Potential difference across "10\Omega=V=IR=2\times 10 =20\ volts"

As the resistors are in parallel, therefore potential difference across all resistors will be same.

"\therefore" Current in "12\Omega=\frac{20}{12} A"

Current in "15\Omega=\frac{20}{15}A"

So, total current "=(2+\frac{20}{12}+\frac{20}{15})A=\frac{120+100+80}{60}=5A"